# Quantum cohomology of projective bundles

### v3

@@ 5.1. Statement
We state a more precise version of Theorem 1.7. Let $\operatorname{QDM}(\mathbb{P}(V))$ be the quantum $D$-module of $\mathbb{P}(V)$ as in §4.1.1. Let $\{\phi_{i}\}_{i=0}^{s}$ be a basis of $H^{*}(B)$ with $\phi_{0}=1$ as before. Then $\{\phi_{i}p^{k}\}_{0\leq i\leq s,0\leq k\leq r-1}$ form a basis of $H^{*}(\mathbb{P}(V))$; we write $\hat{\tau}=\sum_{i=0}^{s}\sum_{k=0}^{r-1}\hat{\tau}^{i,k}\phi_{i}p^{k}$ for the parameter in $H^{*}(\mathbb{P}(V))$. Set
@@
We introduce the localized quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}$ as
$\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}=\operatorname{QDM}(\mathbb{P}(V))\otimes_{\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!].$
It is equipped with the connection $\nabla$ and the pairing $P_{\mathbb{P}(V)}$ as before.
@@
There exist maps $H^{*}(\mathbb{P}(V))\to H^{*}(B)$, $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ with $\varsigma_{j}(\hat{\tau})\in-c_{1}(V)\log\lambda_{j}+H^{*}(B)(\!(q^{-1/r})\!)[\![Q,\hat{\tau}]\!]$, $j=0,\dots,r-1$ with $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$ and an isomorphism
$\Phi\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$
such that
@@
    $\varsigma_{j}(\hat{\tau})+c_{1}(V)\log\lambda_{j}$ is homogeneous of degree $2$ and $\Phi$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}(\mathbb{P}(V))$, $H^{*}(B)$ and the following degree of the variables: $\deg\hat{\tau}^{i,k}=2-\deg\phi_{i}-2k$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$, $\deg q=2r$ and $\deg z=2$;
@@
    $\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}-c_{1}(V)\log\lambda_{j}+O(q^{-1/r})$;
@@
    the Jacobian matrix of $\varsigma(\hat{\tau})=\bigoplus_{j=0}^{r-1}\varsigma_{j}(\hat{\tau})$ along $Q=\hat{\tau}=0$ is $(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})|_{Q=\hat{\tau}=0}=\phi_{i}\lambda_{j}^{k}(1+O(q^{-1/r}))$ and is invertible over $\mathbb{C}(\!(q^{-1/r})\!)$;
@@
    $\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}(\phi_{i}+O(q^{-1/r}))$
where $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}$ is the module
$\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm loc}=H^{*}(B)\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}]\!]$
equipped with the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection of $B$ and the pairing $P_{B}$ (2.8).
@@
Although the maps $\hat{\tau}\mapsto\varsigma_{j}(\hat{\tau})$ contain $\log q$, the pull-back $\varsigma_{j}^{*}\nabla$ of the quantum connection is well-defined by the Divisor Equation. We can eliminate $\log q$ from the statement of the above theorem by considering a coordinate change for the Novikov variables. Consider the map
$\varsigma^{\rm reg}\colon(\hat{\tau},q,Q)\longmapsto(\varsigma,Q_{B})=\left(\{\varsigma_{j}(\hat{\tau})+c_{1}(V)(\log q)/r\}_{j=0}^{r-1},q^{-c_{1}(V)/r}Q\right)$
where $\varsigma$ is a parameter of $\bigoplus_{j=0}^{r-1}H^{*}(B)$ and $Q_{B}$ is the Novikov variable for $B$. The map $\varsigma^{\rm reg}$ does not contain $\log q$. Then the Divisor Equation shows that $\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\nabla$ is the same as $(\varsigma^{\rm reg})^{*}(\nabla^{\oplus r})$. The identification $Q_{B}^{d}=q^{-c_{1}(V)\cdot d/r}Q^{d}$ of the Novikov variables is compatible with their degrees.
@@
As suggested in [28, 24], we expect that the (formal) decomposition of quantum cohomology $D$-modules occurs over formal power series in $z$. We defined the localized quantum $D$-modules over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$, but more precisely we can define them over the ‘homogeneous’ completion $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ consisting of finite sums of homogeneous elements in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$. We observe that $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ is contained in the ring $\mathbb{C}[q^{-1/r^{\prime}},q^{1/r^{\prime}}][\![z]\!]$ of formal power series in $z$. If we have convergence of quantum cohomology with respect to the parameters $q$, $Q$, $\hat{\tau}$, we expect that the decomposition $\Phi$ cannot be analytified in $z$ but should be defined over $\mathcal{O}[\![z]\!]$, with $\mathcal{O}$ the sheaf of analytic functions in $q,Q,\hat{\tau}$. As the picture in [28, 24] suggests, this formal (in $z$) decomposition should be lifted to an analytic decomposition over an angular sector on the $z$-plane, and we expect that the analytic decomposition can be related (at certain values of the parameters) to the semi-orthogonal decomposition [43] of derived categories of coherent sheaves
$D^{b}(\mathbb{P}(V))=\langle D^{b}(B)_{0},\dots,D^{b}(B)_{r-1}\rangle$
via the $\widehat{\Gamma}$-integral structure, where $D^{b}(B)_{i}=D^{b}(B)\otimes\mathcal{O}(i)$.
@@ 5.2. Idea of the proof
[…]
should give a solution for $\operatorname{QDM}(\mathbb{P}(V))$ through the Fourier isomorphism $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\cong\operatorname{QDM}_{S^{1}}(V)$ in Theorem 4.2. Despite the convergence issue of $\Delta_{V}^{\lambda}$ raised in Remark 3.2, we can formally make sense of this Fourier transformation using the stationary phase approximation (i.e. the method of steepest descent). The QRR theorem tells us that this lies in the Givental cone of $B$ and induces projections $\operatorname{QDM}(\mathbb{P}(V))\to\varsigma_{j}^{*}\operatorname{QDM}(B)$, $j=0,\dots,r-1$, where the $r$ choices result from the choice of critical points we use for the stationary phase method.
@@
The QRR operator $\Delta_{V}^{\lambda}$ induces an isomorphism between the quantum $D$-modules of $V$ and $B$ after localizing $\lambda$. The isomorphism is given by a gauge transformation $R(\tau)\in\lambda^{-r/2}\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[z](\!(\lambda^{-1})\!)[\![Q,\tau]\!]$ from $\operatorname{QDM}_{S^{1}}(V)$ to $\sigma^{*}\operatorname{QDM}(B)$ satisfying
@@
The shift operator acts on $\operatorname{QDM}(B)$ by the trivial one $e^{-z\partial_{\lambda}}$, but it acts highly non-trivially on the pull-back by the $\lambda$-dependent coordinate change $\tau\mapsto\sigma(\tau)$. From this point of view, the solution $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$ considered above corresponds to the standard solution $M_{B}(\sigma(\tau))$ for $\sigma^{*}\operatorname{QDM}(B)$.
@@ 5.3. Fourier transformation via stationary phase method
[…]
@@
$\varphi(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})=-r\lambda_{0}+\frac{r}{2}s^{2}+\varphi_{\geq 3}(s,\lambda_{0})\quad\text{with $\varphi_{\geq 3}(s,\lambda_{0})=r\sum_{k=3}^{\infty}\frac{k-1}{k!}\frac{s^{k}}{\sqrt{\lambda_{0}}^{k-2}}$}$
@@
We expand $L(s,\lambda_{0})$ in power series in $s$ and perform the Gaussian integral (5.3) termwise over $s\in\mathbb{R}$ (assuming $z>0$) and arrive at the formal stationary phase asymptotics
[…]
@@
[…]
@@
in terms of the coordinate $s$ around the critical point (given by $\lambda=\lambda_{j}e^{s/\sqrt{\lambda_{j}}}$) and this makes sense in the ring $\mathbb{C}[z,z^{-1},s](\!(\lambda_{j}^{-1/2})\!)$. To show part (3), observe that the map $J\mapsto e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J)$ is homogeneous of degree $-(r-1)$ with respect to the usual grading on $H^{*}_{S^{1}}(V)$ and $\deg q=2r$, $\deg z=2$. This means
@@ 5.4. Fourier transform of the $J$-function
Write $\tau(\lambda)=\sum_{i,k}\tau^{i,k}\lambda^{k}\phi_{i}$ for the parameter in $H^{*}_{S^{1}}(V)$. There exist
$\sigma_{j}$ $\in-c_{1}(V)\log\lambda_{j}+\tau(\lambda_{j})+\mathfrak{m}H^{*}(B)[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
@@
with $\lambda_{j}=q^{1/r}e^{2\pi\sqrt{-1}j/r}$ such that $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$, where
@@
  • $\mathfrak{m}$ ((((resp. $\mathfrak{m}^{\prime}$)))) is the closed ideal of $\mathbb{C}[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ ((((resp. of $\mathbb{C}[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$)))) generated by $q^{-1/r}$ and $q^{-c_{1}(V)\cdot d/r}Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$.
Moreover, $\sigma_{j}+c_{1}(V)\log\lambda_{j}$ and $\upsilon_{j}$ are homogeneous of degree $2$ and $-(r-1)$ respectively with respect to the usual grading on $H^{*}(B)$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$ and $\deg z=2$.
@@
where we set $u=s/\sqrt{\lambda_{j}}$ as before and $Q\to\lambda^{c_{1}(V)}Q$ means $Q^{d}\to\lambda^{c_{1}(V)\cdot d}Q^{d}$. The quantity in the big parenthesis []delimited-[][$\cdots$][ ⋯ ] belongs to $H^{*}(B)[z,z^{-1},s][\![\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and defines a $\mathbb{C}[\![s,\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$-valued point of $\mathcal{L}_{B}$: the Dilaton and String Equations show that $\mathcal{L}_{B}$ is invariant under $e^{-(\frac{r}{2}-1)u}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}$ and the Divisor Equation shows that $\mathcal{L}_{B}$ is invariant under $f(Q)\to e^{-uc_{1}(V)/z}f(Qe^{-uc_{1}(V)})$. By Lemma 2.7, the action of $e^{z(\partial_{s})^{2}/(2r)}$ preserves points on the cone [^7] Strictly speaking, the operator $(z\partial_{s})^{2}/(2r)$ does not vanish at the origin of the formal parameters as required in Lemma 2.7, but the conclusion holds since the action of $e^{z(\partial_{s})^{2}/(2r)}$ is well-defined.. Therefore we have $\tilde{\sigma}_{j}\in H^{*}(B)[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and $\tilde{\upsilon}_{j}\in H^{*}(B)[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ such that $\tilde{\sigma}_{j}\equiv\tau(\lambda_{j})$, $\tilde{\upsilon}_{j}\equiv 1$ modulo the closed ideal in $\mathbb{C}[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ or in $\mathbb{C}[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ generated by $\lambda_{j}^{-1}$ and $Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$ and
@@
As in the proof of Theorem 4.2, $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ can be obtained from the Birkhoff factorization of the matrix formed by the derivatives of the left-hand side with respect to $\tau^{i,0}$. It follows from the homogeneity of $e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))$ that $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ are homogeneous of degree zero after the change $Q\to\lambda_{j}^{-c_{1}(V)}Q$ of variables. By the Divisor Equation again, we obtain $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j})\upsilon_{j}$ with
$\sigma_{j}=-c_{1}(V)\log\lambda_{j}+\tilde{\sigma}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q},\quad\upsilon_{j}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\tilde{\upsilon}_{j}\bigr{|}_{Q\to\lambda_{j}^{-c_{1}(V)}Q}.$

### v4

@@ 5.1. Statement
We now present a more precise version of Theorem 1.7. Let $\{\phi_{i}\}_{i=0}^{s}$ be a basis of $H^{*}(B)$ with $\phi_{0}=1$ as before. Then $\{\phi_{i}p^{k}\}_{0\leq i\leq s,0\leq k\leq r-1}$ form a basis of $H^{*}(\mathbb{P}(V))$. We denote by $\sigma=\sum_{i=0}^{s}\sigma^{i}\phi_{i}\in H^{*}(B)$ the bulk parameter for $B$ and by $\hat{\tau}=\sum_{i=0}^{s}\sum_{k=0}^{r-1}\hat{\tau}^{i,k}\phi_{i}p^{k}\in H^{*}(\mathbb{P}(V))$ the bulk parameter for $\mathbb{P}(V)$. Set
@@
As in Notation 3.1, we use the pair $(q,Q)$ to denote the Novikov variable of $\mathbb{P}(V)$. Recall from §4.1.1 that the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ of $\mathbb{P}(V)$ is defined over the ring $\mathbb{C}[z,q][\![Q]\!]$. Its localized version is defined as
$\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}:=\operatorname{QDM}(\mathbb{P}(V))\otimes_{\mathbb{C}[z,q][\![Q,\hat{\tau}]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!],
equipped with the induced connection $\nabla$ and the pairing $P_{\mathbb{P}(V)}$.
In this section, we introduce an additional copy $Q_{B}$ of the Novikov variable for $B$ and define the quantum $D$-module of $B$ to be the module
$\operatorname{QDM}(B)=H^{*}(B)[z][\![Q_{B},\sigma]\!]$
equipped with the quantum connection $\nabla_{\sigma^{i}},\nabla_{\xi Q_{B}\partial_{Q_{B}}},\nabla_{z\partial_{z}}$ (with $\xi\in H^{2}(B)$) and the $z$-sesquilinear pairing $P_{B}$ (see §2.2). We relate the Novikov variable $Q_{B}$ of $B$ to the variables $(q,Q)$ of $\mathbb{P}(V)$ via the following formula:
(5.2) $Q_{B}^{d}=q^{-c_{1}(V)\cdot d/r}Q^{d}$
for $d\in H_{2}(B,\mathbb{Z})$, which is consistent with the degrees $\deg Q_{B}^{d}=2c_{1}(TB)\cdot d$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$. The formula (5.2) defines an embedding of the coefficient ring $\mathbb{C}[z][\![Q_{B}]\!]$ for $\operatorname{QDM}(B)$ into the ring $\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q]\!]\subset\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q]\!]$. We introduce the following base changes of $\operatorname{QDM}(B)$:
(5.3) $\begin{split}\operatorname{QDM}(B)_{\rm ext}&:=\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![Q_{B},\sigma]\!]}\mathbb{C}[z][\![q^{-1/r^{\prime}},q^{-c_{1}(V)/r}Q,\sigma]\!],\\ \operatorname{QDM}(B)_{\rm ext,loc}&:=\operatorname{QDM}(B)\otimes_{\mathbb{C}[z][\![Q_{B},\sigma]\!]}\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\sigma]\!].\end{split}$
These extended modules are equipped with the quantum connection
$\nabla_{\sigma^{i}}$ $=\nabla_{\sigma^{i}}\otimes 1+1\otimes\partial_{\sigma^{i}}$ $=\partial_{\sigma^{i}}+z^{-1}(\phi_{i}\star_{\sigma})$
$\nabla_{q\partial_{q}}$ $=\nabla_{-(c_{1}(V)/r)Q_{B}\partial_{Q_{B}}}\otimes 1+1\otimes q\partial_{q}$ $=q\partial_{q}-z^{-1}((c_{1}(V)/r)\star_{\sigma})$
$\nabla_{\xi Q\partial_{Q}}$ $=\nabla_{\xi Q_{B}\partial_{Q_{B}}}\otimes 1+1\otimes\xi Q\partial_{Q}$ $=\xi Q\partial_{Q}+z^{-1}(\xi\star_{\sigma})$
$\nabla_{z\partial_{z}}$ $=\nabla_{z\partial_{z}}\otimes 1+1\otimes z\partial_{z}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\sigma})+\mu_{B}$
where $\xi\in H^{2}(B)$, and the $z$-sesquilinear pairing $P_{B}$. These structures are induced from those for $\operatorname{QDM}(B)$ via pull-back along the map $(Q,q)\mapsto Q_{B}$ given by (5.2).
@@
[…]
$\Phi=\bigoplus_{j=0}^{r-1}\Phi_{j}\colon\operatorname{QDM}(\mathbb{P}(V))_{\rm loc}\cong\bigoplus_{j=0}^{r-1}\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$
where $\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$ is the module $H^{*}(B)\otimes\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)[\![Q,\hat{\tau}$]\!] equipped with the pull-back quantum connection $\varsigma_{j}^{*}\nabla$
$\varsigma_{j}^{*}\nabla_{\hat{\tau}^{i,k}}$ $=\partial_{\hat{\tau}^{i,k}}+z^{-1}(\partial_{\hat{\tau}^{i,k}}\varsigma_{j})\star_{\varsigma_{j}(\hat{\tau})}$
[…]
$\varsigma_{j}^{*}\nabla_{z\partial_{z}}$ $=z\partial_{z}-z^{-1}(E_{B}\star_{\varsigma_{j}(\hat{\tau})})+\mu_{B}$
and the pairing $P_{B}$ as in (2.8), such that the following holds with $\lambda_{j}=e^{2\pi\sqrt{-1}j/r}q^{1/r}$:
@@
[…]
@@
    $\varsigma_{j}(\hat{\tau})|_{Q=\hat{\tau}=0}=r\lambda_{j}-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+O(q^{-1/r})$;
@@
[…]
@@
    $\Phi_{j}(\phi_{i}p^{k})|_{Q=\hat{\tau}=0}=\frac{1}{\sqrt{r}}\lambda_{j}^{k-(r-1)/2}(\phi_{i}+O(q^{-1/r}))$.
@@
The identification (5.2) corresponds to a splitting of $\pi_{*}\colon H_{2}(\mathbb{P}(V),\mathbb{Z})\twoheadrightarrow H_{2}(B,\mathbb{Z})$ over $\mathbb{Q}$ defined by the kernel of $c_{1}(T_{\rm vert}\mathbb{P}(V))=rp+\pi^{*}c_{1}(V)$. Here $T_{\rm vert}\mathbb{P}(V)=\operatorname{Ker}(\pi_{*}\colon T\mathbb{P}(V)\to\pi^{*}TB)$ denotes the vertical tangent bundle. This splitting differs from the one in (3.1) defined by the kernel of $p$; however, it is intrinsic to the geometry of the projective bundle $\mathbb{P}(V)\to B$ and is independent of the choice of the vector bundle $V$ (up to tensoring with a line bundle).
@@
We considered the quantum $D$-module $\operatorname{QDM}(\mathbb{P}(V))$ and its decomposition over $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$. More precisely, these can be defined over the ‘homogeneous’ completion $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$, which consists of finite sums of homogeneous elements in $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)$. Because $z$ and $q$ both have positive degrees, $\mathbb{C}[z](\!(q^{-1/r^{\prime}})\!)_{\rm hom}$ is contained in the ring $\mathbb{C}[q^{-1/r^{\prime}},q^{1/r^{\prime}}][\![z]\!]$ of formal power series in $z$. This is relevant to Remark 1.11.
@@ 5.2. Idea of the proof
$s\longmapsto\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)s\,d\lambda$
should give a solution for $\operatorname{QDM}(\mathbb{P}(V))$ through the Fourier isomorphism $\hat{\tau}^{*}\operatorname{QDM}(\mathbb{P}(V))\cong\operatorname{QDM}_{S^{1}}(V)$ in Theorem 4.2. Despite the convergence issue of $\Delta_{V}^{\lambda}$ raised in Remark 3.3, we can formally make sense of this Fourier transformation using the stationary phase approximation (i.e. the method of steepest descent). The QRR theorem tells us that this lies in the Givental cone of $B$ and induces projections $\operatorname{QDM}(\mathbb{P}(V))\to\varsigma_{j}^{*}\operatorname{QDM}(B)_{\rm ext,loc}$, $j=0,\dots,r-1$, where the $r$ choices result from the choice of critical points we use for the stationary phase method.
@@
The QRR operator $\Delta_{V}^{\lambda}$ induces an isomorphism between the quantum $D$-modules of $V$ and $B$ after localizing $\lambda$. The isomorphism is given by a gauge transformation $R(\tau)\in\lambda^{-r/2}\operatorname{End}(H^{*}(B))\otimes\mathbb{C}[z](\!(\lambda^{-1})\!)[\![Q,\tau]\!]$ from $\operatorname{QDM}_{S^{1}}(V)$ to $\sigma^{*}(\operatorname{QDM}(B)|_{Q_{B}\to Q})$ satisfying
@@
The shift operator acts on $\operatorname{QDM}(B)$ by the trivial one $e^{-z\partial_{\lambda}}$, but it acts highly non-trivially on the pull-back by the $\lambda$-dependent coordinate change $\tau\mapsto\sigma(\tau)$. From this point of view, the solution $(\Delta_{V}^{\lambda})^{-1}M_{V}(\tau)$ considered above corresponds to the standard solution $M_{B}(\sigma(\tau))$ for $\sigma^{*}(\operatorname{QDM}(B)|_{Q_{B}\to Q})$.
@@ 5.3. Fourier transformation via stationary phase method
$\int e^{-\varphi(\lambda)/z}\lambda^{-c_{1}(V)/z}\lambda^{-r/2}K(\lambda)d\lambda$
@@
$\varphi(\lambda_{0}e^{s/\sqrt{\lambda_{0}}})=-r\lambda_{0}+\frac{r}{2}s^{2}+\varphi_{\geq 3}(s,\lambda_{0})\quad\text{with $\varphi_{\geq 3}(s,\lambda_{0})=r\sum_{k=3}^{\infty}\frac{k-1}{k!}\frac{s^{k}}{(\sqrt{\lambda_{0}})^{k-2}}$}$
@@
By expanding $L(s,\lambda_{0})$ as a power series in $s$ and performing the Gaussian integral (5.5) termwise over $s\in\mathbb{R}$ (assuming $z>0$), we obtain the formal stationary phase asymptotics
$\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda\sim\sqrt{2\pi z}\cdot e^{r\lambda_{0}/z}\mathscr{F}_{0}(J)$
@@
$zq\partial_{q}\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$ $=\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}\lambda J(\lambda)d\lambda$
$q\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$ $=\int e^{(\lambda+z)\log q/z}(\Delta_{V}^{\lambda})^{-1}J(\lambda)d\lambda$
$=\int e^{\lambda\log q/z}(\Delta_{V}^{\lambda})^{-1}\mathcal{S}J(\lambda)d\lambda$
@@
in terms of the coordinate $s$ around the critical point (given by $\lambda=\lambda_{j}e^{s/\sqrt{\lambda_{j}}}$) and this makes sense in the ring $\mathbb{C}[z,z^{-1},s](\!(\lambda_{j}^{-1/2})\!)$. To show part (3), observe that the map $J\mapsto e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J)$ is homogeneous of degree (r1)-(r-1) with respect to the usual grading on $H^{*}_{S^{1}}(V)$ and $\deg q=2r$, $\deg z=2$. This means
@@ 5.4. Fourier transform of the $J$-function
Write $\tau(\lambda)=\sum_{i,k}\tau^{i,k}\lambda^{k}\phi_{i}$ for the parameter in $H^{*}_{S^{1}}(V)$. Let $M_{B}(\sigma)=M_{B}(\sigma;Q_{B})$ denote the fundamental solution (2.6) of the quantum connection of $B$. There exist
$\sigma_{j}$ $\in-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\tau(\lambda_{j})+\mathfrak{m}H^{*}(B)[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$
@@
with $\lambda_{j}=q^{1/r}e^{2\pi\sqrt{-1}j/r}$ such that $\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=q^{-c_{1}(V)/(rz)}M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$, where
@@
  • $\mathfrak{m}$ ((resp. $\mathfrak{m}^{\prime}$)) is the closed ideal of $\mathbb{C}[\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$ ((resp. of $\mathbb{C}[z][\![q^{-1/r},q^{-c_{1}(V)/r}Q,q^{\bullet/r}\boldsymbol{\tau}]\!]$)) generated by $q^{-1/r}$ and $q^{-c_{1}(V)\cdot d/r}Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$.
Moreover, $\sigma_{j}$ and $\upsilon_{j}$ are homogeneous of degree $2$ and (r1)-(r-1) respectively with respect to the usual grading on $H^{*}(B)$ and the following degree of the variables: $\deg\tau^{i,k}=2-\deg\phi_{i}-2k$, $\deg q=2r$, $\deg Q^{d}=2(c_{1}(TB)+c_{1}(V))\cdot d$ and $\deg z=2$.
@@
where we set $u=s/\sqrt{\lambda_{j}}$ as before and $Q\to\lambda^{c_{1}(V)}Q$ means $Q^{d}\to\lambda^{c_{1}(V)\cdot d}Q^{d}$. The quantity in the big parenthesis $\cdots$] belongs to $H^{*}(B)[z,z^{-1},s][\![\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and defines a $\mathbb{C}[\![s,\lambda_{j}^{-1/2},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$-valued point of $\mathcal{L}_{B}$: the Dilaton and String Equations show that $\mathcal{L}_{B}$ is invariant under $e^{-(\frac{r}{2}-1)u}e^{-\varphi_{\geq 3}(s,\lambda_{j})/z}$ and the Divisor Equation shows that $\mathcal{L}_{B}$ is invariant under $f(Q)\to e^{-uc_{1}(V)/z}f(Qe^{-uc_{1}(V)})$. By Lemma 2.7, the action of $e^{z(\partial_{s})^{2}/(2r)}$ preserves points on the cone [^9] Strictly speaking, the operator $(z\partial_{s})^{2}/(2r)$ does not vanish at the origin of the formal parameters as required in Lemma 2.7, but the conclusion holds since the action of $e^{z(\partial_{s})^{2}/(2r)}$ is well-defined.. Therefore we have $\tilde{\sigma}_{j}\in H^{*}(B)[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ and $\tilde{\upsilon}_{j}\in H^{*}(B)[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ such that $\tilde{\sigma}_{j}\equiv\tau(\lambda_{j})$, $\tilde{\upsilon}_{j}\equiv 1$ modulo the closed ideal in $\mathbb{C}[\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ or in $\mathbb{C}[z][\![\lambda_{j}^{-1},Q,\lambda_{j}^{\bullet}\boldsymbol{\tau}]\!]$ generated by $\lambda_{j}^{-1}$ and $Q^{d}$ with $d\in\operatorname{Eff}(B)\setminus\{0\}$ and
@@
As in the proof of Theorem 4.2, $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ can be obtained from the Birkhoff factorization of the matrix formed by the derivatives of the left-hand side with respect to $\tau^{i,0}$. It follows from the homogeneity of $e^{c_{1}(V)\log\lambda_{j}/z}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))$ that $M_{B}(\tilde{\sigma}_{j})$ and $\tilde{\upsilon}_{j}$ are homogeneous of degree zero after the change $Q\to\lambda_{j}^{-c_{1}(V)}Q$ of variables. By the Divisor Equation again, we obtain $q^{c_{1}(V)/(rz)}\mathscr{F}_{j}(J_{V}(\tau(\lambda)))=M_{B}(\sigma_{j};q^{-c_{1}(V)/r}Q)\upsilon_{j}$ with
$\sigma_{j}=-\frac{2\pi\sqrt{-1}j}{r}c_{1}(V)+\tilde{\sigma}_{j}\bigr|_{Q\to\lambda_{j}^{-c_{1}(V)}Q},\quad\upsilon_{j}=\frac{1}{\sqrt{r}}\lambda_{j}^{-(r-1)/2}\tilde{\upsilon}_{j}\bigr|_{Q\to\lambda_{j}^{-c_{1}(V)}Q}.$